# Private From Whom? Minimal Information Leakage in Auctions

### v1

@@ Abstract
In many auctions, bidders may be reluctant to reveal private information to the auctioneer and other bidders. Among deterministic bilateral communication protocols, reducing what bidders learn requires increasing what the auctioneer learns. A protocol implementing a given social choice rule is on the privacy Pareto frontier if no alternative protocol reveals less to both bidders and the auctioneer. For first-price auctions, the descending protocol and the sealed-bid protocol are both on the privacy Pareto frontier. For second-price auctions, the ascending protocol and the ascending-join protocol of haupt2025contextually are both on the privacy Pareto frontier, but the sealed-bid protocol is not. A designer can flexibly trade off between what bidders learn and what the auctioneer learns by “stitching” different protocols together.
@@ 1 Introduction
In many high-stakes economic settings, unwanted disclosure of agents’ private information can be extremely costly. In the 1960s, the Walt Disney Company extensively used shell corporations to conceal its valuation for the land it purchased to build Walt Disney World.222“[Disney] used aliases and code names, as well as shell and dummy corporations like Project Winter, Project X and Project Future. If word leaked that Walt Disney was buying up huge tracts, prices would shoot sky-high, dooming the project. Eighteen months of secrecy to buy up 27,400 acres.” (Payne 2021). In FCC Spectrum Auction 66, SpectrumCo successfully used early bids to estimate rival bidders’ budgets, allowing SpectrumCo to adjust its auction strategy and save billions of dollars. [^3] In 2006, the Federal Communications Commission conducted Auction 66: Advanced Wireless Services (AWS), netting over 13 billion dollars in revenue. In round nine, “SpectrumCo made a jump bid doubling the prices of the large REAG licenses covering the Northeast and Western United States. This move, which we referred to at the time as the ‘shake-out tactic’ was intended to resolve competitive uncertainty and favorably align relative prices in the auction.” Later on in “round 15, it was possible to forecast that cumulative high bids on the REAG licenses were so high relative to the total budgets in the auction that the smaller licenses would sell for a steep discount. This allowed SpectrumCo, alone among the major bidders, to make an early switch to the smaller licenses.” Such analysis hinged on observing other bidders’ behavior in early rounds: “total winning bids will be forecast early and well by total exposure.” SpectrumCo was able to save over \$1.1 billion to acquire licenses covering the United States, compared to the average price paid by its competitors. (bulow2009winning). As rothkopf1990vickrey and milgrom2020clock describe, bidders may be concerned that their bids reveal their technology level to rivals or reduce the bidder’s leverage in post-auction negotiations; they highlight the importance of privacy in auctions for construction contracts, oil leases, and spectrum. [^4] See rothkopf1990vickrey: “People of our acquaintance with experience in conducting business are reluctant to reveal their true costs or valuations. … A truth-revealing strategy may give away valuable information. It could reveal to potential competitors the extent to which a firm’s technology was superior. Most important, it could reveal to others with whom the firm must subsequently negotiate precisely how much it can yield. … Successful oil lease bidders must deal with drilling contractors, rig owners, and so forth. Successful coal lease bidders must deal with equipment suppliers, railroads, and coal purchasers. Successful construction contract bidders must deal with subcontractors and labor unions.” See milgrom2020clock on spectrum auctions: “The practical significance of UWP [Unconditional Winner Privacy] is that it may alleviate winners’ concerns about misuse of the revealed information, encourage participation by bidders who find it costly to figure out their exact values, and allow the auctioneer to conceal winners’ politically sensitive windfalls.” Faced with these concerns, a designer may wish to choose an auction protocol that preserves bidders’ privacy from both the auctioneer and other bidders. Can any auction format preserve privacy in both dimensions? We identify a stark tradeoff between these objectives: For most choice rules, either the auctioneer or bidders must learn unnecessary information.
We consider a designer, external to the interaction, who constructs a deterministic bilateral communication protocol among an auctioneer and multiple bidders to implement a given choice function. The auctioneer learns information about bidders’ types by sequentially communicating with them, which partitions the space of type profiles. Bidders also learn about other bidders’ types: a bidder’s observed history is only consistent with some realization of other bidders’ types, yielding an analogous partition. Protocol $P$ reveals more to the auctioneer than protocol $P^{\prime}$ if the auctioneer learns a finer partition of the type space under protocol $P$ than protocol $P^{\prime}$. Similarly, protocol $P$ reveals more to bidder $i$ than protocol $P^{\prime}$ if bidder $i$ learns a finer partition under $P$ than $P^{\prime}$.
Under a general condition (the indistinguishable corners condition, a weakening of the corners condition previously seen in the literature), we show that the designer faces a tradeoff between protecting bidders’ privacy from one another or from the auctioneer. Intuitively, to protect bidders’ privacy from the auctioneer, the protocol must refrain from asking bidders questions about their type when doing so is unnecessary to compute the outcome. However, which questions are necessary depends on what the auctioneer knows so far; by conditioning such questions on other bidders’ types, bidders infer what the auctioneer has learned and, hence, other bidders’ private information.
We then use this framework to compare the privacy tradeoffs between different protocols. Protocol $P$ privacy dominates protocol $P^{\prime}$ if protocol $P$ reveals less to both the auctioneer and other bidders than protocol $P^{\prime}$ does, with at least one of these holding strictly. Under these orderings, a protocol lies on the privacy Pareto frontier if no protocol that implements the same underlying choice rule privacy dominates it. We then apply our results to first-price and second-price auctions.
In a first-price auction setting, the sealed-bid first-price auction reveals more information to the auctioneer than the descending auction, but reveals less information to bidders. In a sealed-bid first-price auction, the auctioneer learns every bidder’s value, but each bidder learns only whether he won or lost. In a descending auction, the auctioneer learns less information by no longer querying any bidder once some bidder claims the good. However, each losing bidder now learns the winning bidder’s value based on the final question the auctioneer asks them before the auction ends. We show that both protocols lie on the privacy Pareto frontier.
In a second-price auction setting, the ascending-join protocol, proposed by haupt2025contextually, is one protocol that reveals minimal information to the auctioneer. However, other protocols reveal less information to bidders. In particular, the ascending-join protocol does not privacy-dominate the standard ascending protocol. This complicates the privacy rationale for using the ascending-join protocol: a privacy-conscious designer may prefer the ascending protocol. Privacy dominance still rules out some protocols: in particular, the ascending auction privacy dominates the sealed-bid second-price auction. We then quantify violations of the indistinguishable corners condition to show that the ascending auction lies on the privacy Pareto frontier.
@@
Our results highlight that the designer often must trade off between revealing minimal information to the auctioneer and revealing minimal information to bidders. The designer’s preferred choice may depend on the specific interaction taking place. In some settings, such as a one-time auction for a procurement contract, preventing bidders from learning about rival bidders’ technological capabilities may be more important. In other settings, such as repeated online ad auctions (in which the auctioneer may set future reserve prices based on bidders’ values), preserving privacy from the auctioneer may be more important. As such, privacy concerns alone may not distinguish among the auction formats that lie on the frontier: protocols that initially seem to be “more private” may actually release more private information to participants. The designer may also value other criteria, such as preferring a static protocol (which requires less time to run) or an obviously strategyproof protocol (which requires less reasoning from participants). We do not take a stand on balancing these criteria with privacy concerns.
@@
Our framework restricts the usage of alternative tools such as cryptography or an external mediator. [^5] Formally, our designer does not rely on a trusted randomization device, as in izmalkov2011perfect, we make no computational hardness assumptions, and assume that the designer also chooses bidder strategies leading to them being invertible. Alternative mechanisms come with their own drawbacks. [^6] Participants must understand or trust that a cryptographic scheme works as intended for the mechanism to be successful. Many alternative mechanisms also require players to communicate in a different manner, such as through an underlying blockchain or by encoding messages in specific ways themselves. Relying on a separate auctioneer or platform requires bidders to trust that the auction host will not divulge any information that they learn. Finally, alternative tools may not exist in certain settings (many auctions are run in certain ways due to historical significance), may become obsolete (for example, if quantum computers are able to break cryptographic schemes), or are too costly for other reasons (in spectrum auctions, where a graph coloring problem must be solved quickly in every iteration, any increase in time complexity could be prohibitive). With that in mind, our results provide a calculus for deciding when the costs of moving to an alternative scheme are worthwhile. A designer may first wish to understand what privacy protections are feasible in the absence of cryptography before deciding whether to use cryptographic protocols; our work provides this characterization.
@@
The remainder of the paper proceeds as follows. Section˜2 reviews related literature. Section˜3 introduces our model, presents our notions of bidder and auctioneer privacy, and comments on the incentive properties of protocols. Section˜4 contains our main characterizations: an impossibility theorem identifying the tradeoff between bidders and the auctioneer learning unnecessary information, and a stitching lemma demonstrating that the tradeoff is flexible. Section˜5 and Section˜6 apply our results to first and second-price auctions respectively, situating common auction formats relative to the frontier. Section˜7 concludes.
@@ 2 Related Literature
Our work is related to the notions of unconditional privacy (which has the strict requirement of zero privacy violations) and contextual privacy (which only analyzes privacy from the auctioneer). brandt2008privacy consider implementing auctions via bilateral communication among bidders without revealing any information, a condition known as unconditional privacy (and considered earlier by chor1991boolean). We instead consider settings with an auctioneer and characterize privacy tradeoffs even when unconditional privacy is unattainable. milgrom2020clock and ausubel2004auction consider the related notion of unconditional winner privacy, which reveals no unnecessary information about the winner’s value; in contrast, we consider the privacy of all bidders. haupt2025contextually characterizes protocols that implement choice rules while revealing minimal unnecessary information to the auctioneer, which they call contextual privacy. We show that contextual privacy typically conflicts with the desire to reveal minimal unnecessary information to bidders. In turn, protocols which are not maximally contextually private may be better at protecting bidders’ privacy, leading us to consider a distinct privacy Pareto frontier. Finally, Liu2025WorstCasePrivacy quantifies the worst-case privacy risk of information structures and studies optimal mechanisms subject to a “budget” of privacy violations. We instead study the maximally privacy-preserving way to implement a given allocation rule. Additionally, Liu2025WorstCasePrivacy considers inferences an outside observer can make in non-auction settings, which differs from information leakages to the bidders and the auctioneer (as they gain additional information by participating in the auction itself) and are a larger concern in auctions.
@@
A broader literature considers protocols that use trusted randomization or computational hardness assumptions. izmalkov2011perfect provide a way to implement any choice rule in the presence of a trusted randomization device that fully preserves privacy. [^7] The implementation in izmalkov2011perfect is, informally speaking, somewhat complex. This may be another motivation for the designer to choose protocols that do not rely on trusted randomization. In contrast, we restrict attention to protocols that do not require trusted randomization, and we study the tradeoffs between protocols commonly used in practice. This provides a benchmark for what privacy protections are possible without resorting to cryptographic protocols. See alvarez2020survey for a survey of cryptographic protocols in auctions and see bogetoft2009sugarbeets for applications to an agricultural auction.
@@
Other literature discusses the importance of bidders’ privacy in real-world settings, typically nesting a mechanism in a larger model. rothkopf1990vickrey consider bidders who are reluctant to reveal their true valuations in Vickrey auctions due to post-auction bargaining. milgrom2020clock, ausubel2004auction and mcmillan1994spectrum discuss the importance of protecting winner privacy in spectrum auctions and multi-unit auctions. dworczak2020aftermarkets models an aftermarket which is affected by the information revealed in an auction. We are motivated by similar concerns as this literature, and we provide a framework to compare whose privacy is compromised by a given protocol, and who their information is revealed to.
@@
Finally, our work relates to the literature on mechanism design with partial commitment. AkbarpourLi2020CredibleAuctions and komo2025shill consider auctioneer deviations from a specified mechanism: Whether such deviations are profitable are informed by what the auctioneer learns while whether such deviations are detectable are disciplined by what bidders learn. grigoryan2023auditability propose an index to measure the extent to which agents can detect deviations from the mechanism. In contrast to this literature, we consider an auctioneer who faithfully executes a specified protocol: Privacy concerns persist even if the auctioneer can commit to the protocol used.
@@ 3 Model
3.1 Primitives
Our model resembles that of haupt2025contextually and the messaging game of AkbarpourLi2020CredibleAuctions. To analyze bidder privacy, we utilize the notion of an bidder’s experience developed by AkbarpourLi2020CredibleAuctions. Our characterizations are type profile by type profile and abstract away from (Bayesian) incentive constraints, so we do not need to work with beliefs. The model consists of the following:
  1. 1.
    A finite set of bidders $N=\{1,2,...,n\}$ with generic element $i$.
  2. 2.
    An auctioneer $A$ that communicates with the bidders and chooses the final outcome.
  3. 3.
    A finite type space $\bm{\Theta}=\times_{i\in N}\Theta_{i}$.
  4. 4.
    A set of outcomes $X$.
  5. 5.
    Bidder utilities $u_{i}:X\times\Theta_{i}\to\mathbb{R}$.
  6. 6.
    A partition $\Omega_{i}$ of $X$, describing which outcomes bidder $i$ can inherently distinguish. [^8] In auction settings, we assume that a bidder observes their allocation and payment at the conclusion of the auction. Suppose $\{(q_{i},t_{i})\}_{i=1}^{N}\in X$ represents an outcome with allocation $q_{i}\in\{0,1\}$ and transfer $t_{i}$ for bidder $i$. Then $\{(q_{i},t_{i})\}_{i=1}^{N}\sim_{\Omega_{i}}\{(q_{i}^{\prime},t_{i}^{\prime})\}_{i=1}^{N}$ if $q_{i}=q_{i}^{\prime}$ and $t_{i}=t_{i}^{\prime}$. For example, bidder $i$ cannot inherently distinguish between two different outcomes which both result in $i$ losing the auction.
  7. 7.
    A social choice rule $\phi:\bm{\Theta}\to X$ that the designer would like to implement.
We consider a fixed choice rule $\phi$, such as a second-price auction rule or a first-price auction rule. The designer may account for many tradeoffs in choosing $\phi$ (such as efficiency, revenue, or desirable incentive properties). For a given choice rule $\phi$, the designer still has considerable latitude to choose an extensive-form implementation of $\phi$; we describe this frontier of implementing the chosen choice rule.
We model the designer as external to the interaction between bidders and the auctioneer. In many mechanism design settings, the economist that advises organizations on what mechanism to use is not a participant themselves, and simply seeks to identify tradeoffs, bounds on what is inherently feasible, and how to best achieve certain objectives. To implement $\phi$, the designer chooses an extensive-form game $G=(H,i,\{\mathcal{I}\}_{i\in N\cup\{A\}},\{\mathcal{A}_{i}\}_{i\in N\cup\{A\}},\mathcal{A},I)$ of finite length and perfect recall with notation specified as follows:
@@ 3.1 Primitives
Following AkbarpourLi2020CredibleAuctions, a game $G$ generates a bilateral communication game for the auctioneer to implement the choice function by repeatedly querying successive bidders and observing their actions before moving to successor information sets until choosing an outcome at a terminal history. [^9] We can formally write down this bilateral communication game as follows: Let the auctioneer’s message space be the union of all players’ information sets, $\cup_{i}\mathcal{I}_{i}$. Then at each stage, according to $G$, 1. The auctioneer either selects an outcome (finishing the game) or proceeds to step 2. 2. The auctioneer sends message $I_{i}\in\mathcal{I}_{i}$ to bidder $i$. 3. Bidder $i$ observes $I_{i}$ and chooses action $a\in A(I_{i})$. 4. The auctioneer observes $a$. 5. Return to step 1. This setup embeds an assumption that the auctioneer cannot “commit to forget” the content of a bidder’s message.
We limit our attention to deterministic strategy profiles. An interim strategy for bidder $i$ is a map $\sigma_{i}:\mathcal{I}_{i}\to A_{i}$ satisfying $\sigma_{i}(I_{i})\in A(I_{i})$. Let $\Sigma_{i}$ denote the space of interim strategies for bidder $i$. An ex-ante strategy for bidder $i$ is a map $S_{i}:\Theta_{i}\to\Sigma_{i}$ mapping from types to interim strategies. A protocol $P$ is a pair $P=(G,\bm{S})$ where $\bm{S}=\{S_{i}\}_{i\in N}$ is some fixed profile of deterministic strategies for $G$.
@@
Each type profile $\bm{\theta}$ yields a sequence of information sets and bidder actions at those information sets. Let $I_{i}^{k}(\bm{\theta})$ denote the $k$’th information set reached at which bidder $i$ is called to play, and let $a_{i}^{k}(\bm{\theta})$ denote their action chosen at this information set. Let $I^{k}(\bm{\theta})$ denote the $k$’th information set reached in the game, and let $a^{k}(\bm{\theta})$ denote the action chosen at this information set. Let $z(\{\sigma_{i}\})$ denote the terminal node reached via playing interim strategies $\{\sigma_{i}\}$ so $z(\bm{S}(\bm{(}\theta))$ is the terminal reached via playing ex-ante strategies $\bm{S}$ at type profile $\bm{\theta}$; protocol $P$ implements social choice rule $\phi$ if for all $\bm{\theta}$, $\phi(\bm{\theta})=g(z(\bm{S}(\bm{\theta}))).$

### v2

@@ Abstract
In many auctions, bidders may be reluctant to reveal private information to the auctioneer and other bidders. Among deterministic bilateral communication protocols, reducing what bidders learn requires increasing what the auctioneer learns. A protocol implementing a given social choice rule is on the Privacy Frontier if no alternative protocol reveals less to both bidders and the auctioneer. For first-price auctions, the descending protocol and the sealed-bid protocol are both on the Privacy Frontier. For second-price auctions, the ascending protocol and the ascending-join protocol of Haupt and Hitzig (2025) are both on the Privacy Frontier, but the sealed-bid protocol is not. We provide sufficient conditions for a protocol to be on the Privacy Frontier and devise alternative protocols on the Privacy Frontier for first-price auctions that allow the designer to flexibly trade off between privacy from bidders and the auctioneer.
@@ 1 Introduction
Bidders participating in auctions may value keeping their information private from both the auctioneer and other bidders. For example, bidders may be concerned that their bids reveal their technology level to rivals or reduce their leverage in post-auction negotiations. [^2] As Rothkopf et al. (1990) describe, “People of our acquaintance with experience in conducting business are reluctant to reveal their true costs or valuations. … It could reveal to potential competitors the extent to which a firm’s technology was superior. Most important, it could reveal to others with whom the firm must subsequently negotiate precisely how much it can yield. … Successful oil lease bidders must deal with drilling contractors, rig owners, and so forth. Successful coal lease bidders must deal with equipment suppliers, railroads, and coal purchasers. Successful construction contract bidders must deal with subcontractors and labor unions.” In the context of spectrum auctions, Milgrom and Segal (2020) write, “The practical significance of UWP [Unconditional Winner Privacy] is that it may alleviate winners’ concerns about misuse of the revealed information, encourage participation by bidders who find it costly to figure out their exact bid, and allow the auctioneer to conceal winners’ politically sensitive windfalls.” Protecting bidders’ private information may encourage bidder participation and prevent the auctioneer from learning sensitive information. [^3] McMillan (1994) describes a New Zealand spectrum auction in which the auction revealed a large gap between the winner’s willingness to pay and her actual payment, resulting in public criticism of the government. A government auctioneer may wish to avoid criticism by reducing the amount she learns about bidders’ values. To what extent can different auction formats protect bidders’ private information from both the auctioneer and other bidders?
Consider a designer choosing between a sealed-bid first-price auction and a descending auction. In a sealed-bid first-price auction, the auctioneer observes every bidder’s exact bid, including losing bids. A descending auction seems to better protect bidders’ privacy: the auction ends when the highest bidder claims the item, so the auctioneer no longer observes exact losing bids. However, a descending auction reveals more information to bidders: each losing bidder now infers that the winning bid is exactly the price at which the auction stops. In contrast, in a sealed-bid first-price auction bidders only learn whether they won or lost.
What drives this tradeoff? Suppose an auctioneer is running a descending auction and a bidder claims the object. The outcome is already determined, so further questions from the auctioneer are unnecessary. But if the auctioneer stops asking questions to some bidder, that bidder infers that a competitor won the auction at the last price they were offered. In general, protecting bidders’ privacy from the auctioneer requires the auctioneer to not ask bidders questions about their type when the outcome is identical regardless of their answer. However, whether a question is necessary depends on what the auctioneer already knows about other bidders’ types: By conditioning which questions to ask on their current information, the auctioneer reveals other bidders’ private information.
This stark tradeoff holds more generally. Consider a designer implementing a given choice rule, which maps from a type space to a set of possible outcomes. [^4] In the single-unit auction setting, a choice rule describes the allocation of the good and agents’ transfers as a function of bidder’ types. The first-price auction choice rule and second-price auction choice rule are two examples. We focus on the single-unit auction setting, but our model and impossibility theorem apply to arbitrary choice rules. To implement the choice rule, the designer chooses a deterministic bilateral communication protocol among an auctioneer and bidders. Under a general condition on the choice rule, implementation necessitates either (1) the auctioneer learning information about every bidder that is not pivotal to the outcome or (2) bidders learning about other bidders’ types beyond what the outcome alone reveals. Consequently, the designer must choose how information leakages are distributed between the auctioneer and bidders. We analyze this understudied additional dimension of privacy—bidders learning about other bidders’ private information.
Given this tradeoff, some privacy violations are inevitable. However, some protocols can still be improved upon from a privacy perspective. Consider now a designer choosing between a sealed-bid second-price auction and an ascending auction. In a sealed-bid second-price auction, the auctioneer learns all bidders’ values; losing bidders learn that the winning bidder’s value is higher than their own; and the winning bidder learns the second-highest value via the price they pay. In an ascending auction, the auctioneer no longer learns the winning bidder’s valuation (the auction ends when the second-to-last bidder drops out), but what bidders learn remains unchanged. [^5] Losing bidders still learn only that the winning bidder’s value is higher than their own; the winning bidder still learns the second-highest value via the price they pay. The ascending auction privacy dominates the sealed-bid second-price auction: It reveals strictly less information to the auctioneer and the same information to bidders. In general, a protocol lies on the Privacy Frontier if it is not privacy dominated by any protocol implementing the same underlying choice rule.
We apply our framework to study the implementation of first-price auctions and second-price auctions. For the first-price auction choice rule, both the sealed-bid first-price auction and the descending auction lie on the Privacy Frontier. We then propose two new protocols on the Privacy Frontier. The descending then sealed-bid protocol trades off privacy from bidders and the auctioneer at different ranges of bidder values. The descending-holdout protocol protects one selected bidder’s privacy from both bidders and the auctioneer.
To implement the second-price auction choice rule, the ascending-join protocol, proposed by Haupt and Hitzig (2025), reveals minimal information to the auctioneer. [^6] Intuitively, the ascending-join protocol orders the bidders and runs an ascending auction with two bidders at a time, until one bidder drops out, then brings the next bidder into the ascending auction. This prevents the auctioneer from learning the value of later bidders, if later bidders have low values and would lose the auction. We elaborate on this protocol in Section 6. However, other protocols reveal less information to bidders. In particular, the ascending-join protocol does not privacy dominate the standard ascending protocol, which reveals minimal information to bidders. We show both the ascending and the ascending-join auctions are on the privacy frontier, but the sealed-bid second-price auction is not.
We then generalize our techniques by providing sufficient conditions to determine whether a given protocol to lie on the Privacy Frontier. These conditions allow a designer to evaluate a protocol only using model primitives and a description of what information the protocol reveals. Our sufficient conditions apply to any type space and choice rule, enabling analysis beyond the structure of auction settings.
@@
Our results give a designer tools to balance revealing information to the auctioneer with revealing information to bidders. The designer’s preferred choice may depend on the specific interaction taking place, which we do not take a stand on. In some settings, such as a one-time auction for a procurement contract, preventing bidders from learning about rival bidders’ technological capabilities may be more important. In other settings, such as repeated auctions for a good (in which the auctioneer may set future reserve prices based on bidders’ values), preserving privacy from the auctioneer may be more important. [^7] Such settings could include repeated auctions for online ads or natural resources.
@@
Our framework of deterministic bilateral communication restricts the usage of alternative tools such as cryptography or an external mediator. [^8] Formally, our designer does not rely on a trusted randomization device as in Izmalkov et al. (2011), and we make no computational hardness assumptions. These alternative tools may be costly to implement and come with their own limitations. Participants must understand or trust that a cryptographic scheme works as intended for the mechanism to be successful. Relying on an external mediator requires bidders to trust that the mediator will not divulge any information. Finally, alternative tools may be inaccessible, become obsolete, or be prohibitively costly for alternative reasons. [^9] In many auctions in practice, the auctioneer and bidders communicate directly without a mediator. These include art and antique auctions (at Sotheby’s or Christie’s), and some auctions for resources such as timber and natural gas. Developments in quantum computing (Cherney (2025)) may undermine the security guarantees of cryptographic schemes. In spectrum auctions, where a graph coloring problem must be solved quickly in every iteration, any increase in time complexity could be prohibitive. With that in mind, our results provide a calculus for deciding when the costs of moving to an alternative scheme are worthwhile. A designer may first wish to understand what privacy protections are feasible in the absence of cryptography before deciding whether to use cryptographic protocols; our work provides this characterization.
@@
The remainder of the paper proceeds as follows. Section˜2 reviews related literature. Section˜3 introduces our model. Section˜4 defines our notion of privacy and states our impossibility theorem, identifying the tradeoff between revealing information to bidders and the auctioneer. Section˜5 and Section˜6 apply our results to first and second-price auctions respectively, situating common auction formats relative to the frontier. Section˜7 provides sufficient conditions for a protocol to lie on the frontier. Section˜8 concludes. All omitted proofs are in Appendix B: Omitted Proofs.
@@ 2 Related Literature
Our work is related to the notions of unconditional privacy (which has the strict requirement of zero privacy violations) and contextual privacy (which only analyzes privacy from the auctioneer). Brandt and Sandholm (2008) consider implementing auctions via bilateral communication among bidders without revealing any information, a condition known as unconditional privacy (and considered earlier by Chor and Kushilevitz (1991)). We instead consider settings with an auctioneer and characterize a privacy frontier even when unconditional privacy is unattainable. Milgrom and Segal (2020) and Ausubel (2004) consider the related notion of unconditional winner privacy, which reveals no unnecessary information about the winner’s value; in contrast, we consider the privacy of all bidders. Privacy from the auctioneer is related to relative informativeness (Mackenzie and Zhou (2022), Segal (2007)), but we further analyze what bidders learn from the mechanism. Haupt and Hitzig (2025) characterizes protocols that implement choice rules while revealing minimal information to the auctioneer, which they call contextual privacy. We show that contextual privacy typically conflicts with the desire to reveal minimal information to bidders, leading to a distinct Privacy Frontier.
Other approaches to privacy include Liu (2025), which quantifies the worst-case privacy risk of information structures and studies optimal mechanisms subject to a “budget” of privacy violations. We instead study the maximally privacy-preserving implementation of a given allocation rule. Häfner et al. (2025) models agents who may learn other agents’ actions through channels outside of the mechanism; we study what information must be revealed by the mechanism itself. In contrast with Eilat et al. (2021), who propose a quantitative measure of information that a mechanism reveals to the designer, we study what agents learn about each other during the mechanism. The literature on differential privacy (Dwork (2006), Pai and Roth (2013)) focuses on what a designer learns from a mechanism’s outcome; we broaden the focus to what both the auctioneer and agents learn from playing the mechanism itself.
@@
A broader literature considers protocols that use trusted randomization or computational hardness assumptions. Izmalkov et al. (2011) provide a way to implement any choice rule in the presence of a trusted randomization device that fully preserves privacy. See Alvarez and Nojoumian (2020) for a survey of cryptographic protocols in auctions and see Bogetoft et al. (2009) for applications to an agricultural auction. In contrast, we provide a benchmark for what privacy protections are possible without resorting to cryptographic protocols or trusted randomization.
@@
Other literature discusses the importance of bidders’ privacy in real-world settings, typically nesting a mechanism in a larger model. Rothkopf et al. (1990) consider bidders who are reluctant to reveal their true valuations in Vickrey auctions due to post-auction bargaining. Milgrom and Segal (2020), Ausubel (2004) and McMillan (1994) discuss the importance of protecting winner privacy in spectrum auctions and multi-unit auctions. Dworczak (2020) models an aftermarket which is affected by the information revealed in an auction. We are motivated by similar concerns as this literature, and we provide a framework to compare whose privacy is compromised by a given protocol, and who their information is revealed to.
@@
Finally, our work relates to the broader literature on extensive-form mechanism design. Segal (2013) utilizes extensive-form protocols to study the communication complexity of verifying that a given allocation rule is faithfully executed. Akbarpour and Li (2020), Grigoryan and Möller (2023), Banchio et al. (2025), and Komo et al. (2025) consider an auctioneer or mediator deviating from a specified mechanism. In contrast to this literature, we consider an auctioneer who faithfully executes a specified protocol. Privacy concerns persist even if the auctioneer can commit to the protocol used. Privacy may also be a component of the credibility tradeoff: Whether such deviations are profitable are informed by what the auctioneer learns, while whether such deviations are detectable are disciplined by what bidders learn.
@@ 3 Model
Consider a finite set of agents (bidders) $N=\{1,2,...,n\}$, with generic element $i$. Bidder $i$ draws values from type space $\Theta_{i}$, independently of other bidders, and $\bm{\Theta}=\times_{i\in N}\Theta_{i}$. $X$ denotes the set of possible outcomes. Bidder $i$ has utility function $u_{i}:X\times\Theta_{i}\to\mathbb{R}$.
The designer seeks to implement a fixed choice rule $\phi:\bm{\Theta}\to X$. In the auction setting, this rule describes agents’ transfers and allocation as a function of the type profile. Examples include a second-price auction rule or a first-price auction rule. The designer may consider other tradeoffs (such as efficiency, revenue, or desirable incentive properties) when choosing $\phi$. The choice rule may also simply be exogenously imposed for institutional reasons.
For a given choice rule $\phi$, the designer still has considerable flexibility to choose an extensive-form implementation of $\phi$; privacy violations stem from the particular extensive-form implementation chosen. Formally, the designer chooses an extensive-form game $G=(H,i,\{\mathcal{I}_{i}\}_{i\in N},\{A_{i}\}_{i\in N},A,I,g)$ of finite length and perfect recall with notation specified in Table˜1.101010We model the designer as external to the interaction between bidders and the auctioneer. In many settings, the designer that advises organizations on what mechanism to use is not a participant themselves, and simply seeks to identify tradeoffs, bounds on what is inherently feasible, and how to best achieve certain objectives.
@@
We limit our attention to deterministic strategy profiles and abstract away from incentives. [^11] Appendix A: Pruning and Incentive Compatibility discusses pruning protocols and establishes that every protocol implementing a dominant strategy incentive compatible choice rule is ex-post incentive compatible. Restricting to deterministic strategies is without loss due to the requirement that a protocol must implement the desired choice rule (with probability one): If two types take the same sequence of actions with positive probability yet the choice rule is still implemented, then these two types never needed to be distinguished. Thus, instead of playing a mixed strategy, the two types could have been pooled, leading to even less information leaked to bidders or the auctioneer than mixing. As such, our results can also be interpreted as a tension between privacy and implementation: If privacy is a sufficiently large concern and bidders rely on randomization, then no deterministic choice rule is implementable. An interim strategy for bidder $i$ is a map $\sigma_{i}:\mathcal{I}_{i}\to A_{i}$ satisfying $\sigma_{i}(I_{i})\in A(I_{i})$ for all $I_{i}$. Let $\Sigma_{i}$ denote the space of interim strategies for bidder $i$. An ex-ante strategy for bidder $i$ is a map $S_{i}:\Theta_{i}\to\Sigma_{i}$ from types to interim strategies. A protocol $P$ is a pair $P=(G,\bm{S})$ where $\bm{S}=\{S_{i}\}_{i\in N}$ is a fixed profile of ex-ante strategies for $G$.
@@
Each type profile $\bm{\theta}$ yields a sequence of information sets and bidder actions at those information sets. Let $I_{i}^{k}(\bm{\theta})$ denote the $k$th information set reached at which bidder $i$ is called to play, and let $a_{i}^{k}(\bm{\theta})$ denote their action chosen at this information set. Let $I^{k}(\bm{\theta})$ denote the $k$th information set reached in the game, and let $a^{k}(\bm{\theta})$ denote the action chosen at this information set. Let $z(\{\sigma_{i}\})_{i\in\mathcal{I}}$ denote the terminal node reached via playing interim strategies $\{\sigma_{i}\}_{i\in\mathcal{I}}$ so $z(\bm{S}(\bm{\theta}))$ is the terminal reached via playing ex-ante strategies $\bm{S}$ at type profile $\bm{\theta}$. Protocol $P$ implements social choice rule $\phi$ if for all $\bm{\theta}$, $\phi(\bm{\theta})=g(z(\bm{S}(\bm{\theta}))).$
We do not restrict the incentive properties of $P=(G,\bm{S})$, which only strengthens our two classes of results. Requiring $P$ to satisfy a particular equilibrium concept only strengthens our impossibility result. When we show protocols lie on the Privacy Frontier, we separately check that they also satisfy desirable incentive properties. [^12] In the second-price auction setting, the given protocols are ex-post incentive compatible or dominant-strategy incentive compatible. In the first-price auction setting, the given protocols form a perfect Bayesian equilibrium. Protocols that lie on the Privacy Frontier, by definition, are not privacy dominated by any protocol: they are therefore not dominated by any protocol satisfying desired incentive properties.